OpenStudy (anonymous):

You need a 90% alcohol solution. On hand, you have a 60 mL of a 20% alcohol mixture. You also have 95% alcohol mixture. How much of the 95% mixture will you need to add to obtain the desired solution?

OpenStudy (anonymous):

Let x and y be the amount 90% mixture and 95% mixture respectively. Solve the following equations for x and y.\[\{0.20*60+0.95x=0.90y,y=x+60\} \]\[\{x=840 \text{ ml},y=900 \text{ ml}\} \]